The accounting records of Old Towne Electronics show the following data.
Beginning inventory 3,110 units at $5
Purchases 7,800 units at $8
Sales 9,060 units at $10
Determine cost of goods sold during the period under a periodic inventory system using the FIFO method. (Round answer to 0 decimal places, e.g. 1,250.)

FIFO

Cost of goods sold during the period $

Determine cost of goods sold during the period under a periodic inventory system using the LIFO method. (Round answer to 0 decimal places, e.g. 1,250.)

LIFO

Cost of goods sold during the period $

Determine cost of goods sold during the period under a periodic inventory system using the average-cost method. (Round average-cost per unit to 3 decimal places, e.g. 12.502 and final answer to 0 decimal places, e.g. 1,250.)

Average-cost

Cost of goods sold during the period $
Answer #1

1.) FIFO Method :- Purchases made earlier will be sold first.

Cost of goods sold = (3,110 x $5) + (5,950 x $8)

= 15,550 + 47,600 = $63,150

2.) LIFO Method :- Purchases made at last will be sold first.

  Cost of goods sold = (7,800 x $8) + (1,260 x $5)

= 62,400 + 6,300 = $68,700

3.) Average cost method :-

Average cost per unit = [(3,110 x 5) + (7,800 x 8)]/(3,110 + 7,800)

= $7.145 per unit

   Cost of goods sold = 9,060 x $7.145 = $64,734